HUBBLE SPACE TELESCOPE DAILY REPORT #2368 PERIOD COVERED: 0000Z (UTC) 05/07/99 - 0000Z (UTC) 05/10/99 Daily Status Report as of 130/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 7949 (CCD Dark and Bias Monitor -- Cont. to June 99) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a CCD dark calibration and a bias monitoring program. It was also scheduled to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.3 Completed WF/PC-2 7279 (The Impact of Cluster Substructure on Galaxies: A Case Study in ABELL 2125) The WF/PC-2 was used to make observations of the Abell-2125 galactic cluster (a.k.a., the Butcher-Oemler Cluster) in order to study the impact of the substructure of the cluster on galaxies. The VLA data show that, as well as a large blue-galaxy fraction {0.19} for its redshift, it has an unprecedented number of radio galaxies {30 with confirmed redshifts}, evenly split between old stellar populations and starbursting galaxies. As described in the previous report and in HSTAR 7122, the acquisition for this proposal defaulted to fine lock back-up on one FGS only, possibly affecting four observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.4 Completed STIS/CCD 8393 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation IV) The Space Telescope Imaging Spectrograph (CCD) was used to make POMS, non-scripted, low galactic latitude observations made in parallel with the prime science observations. As described in the previous report and in HSTAR 7122, the acquisition for the first iteration of this proposal defaulted to fine lock back-up on one FGS only, possibly affecting four observations. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.5 Completed STIS/CCD 8080 (STIS Spectroscopic CCD Flatfielding -- Revised) The Space Telescope Imaging Spectrograph (CCD) was used to make CCD flat fielding calibration observations in the spectroscopic mode. The observations were completed as planned, and no anomalies were reported. 1.6 Completed WF/PC-2 6359 (Core Properties of the Bulges of Spiral Galaxies) The WF/PC-2 was used to make snapshot survey observations of spiral galaxies ESO323G58 and ESO502G20 in order to study their central bulges. The aim is to investigate the core properties, correlations with global galactic properties, and the frequency of occurrence of multiple and/or miscentered nuclei. The observations were completed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD 8152 (Mineralogy and Weathering History of Mars) The Space Telescope Imaging Spectrograph (CCD) was used to make high spatial resolution and moderate spectral resolution 5500 Angstrom to 10,000 Angstrom STIS long-slit imaging spectroscopy observations of the planet Mars. This is to identify and globally map the planet's iron-bearing surface mineralogy. As described in the previous report and in HSTAR 7124, the acquisition for this proposal defaulted to fine lock back-up on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. ********************************************************* Upon completion of the above proposal, the inertial hold safemode event occurred. The following scheduled proposals did not execute: A. Two Sets of STIS/CCD 7949 (CCD Dark and Bias Monitor -- Cont. to June 99) B. Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) C. STIS/CCD Eleven Sets of 8393 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation IV) D. STIS/CCD 8080 (STIS Spectroscopic CCD Flatfielding -- Revised) E. WF/PC-2 6359 (Core Properties of the Bulges of Spiral Galaxies) F. WF/PC-2 7504 (Calibration of Nearby Type Ia Supernovae as Standard Candles: The Next Three-Year Step) G. Four Sets of WF/PC-2 8053 (WF/PC-2 Cycle-7 Earth Flats -- Part II) H. WF/PC-2/STIS/CCD 7615 (Determination of the Mass Function of a Globular Cluster Through Gravitational Microlensing) I. Four Sets of WF/PC-2 6512 (The Formation of the Galaxy: Completing the Age Profile of the Outer Halo Globular Clusters) J. Two Sets of WF/PC-2 6362 (Supernova Host Galaxy) K. Two Sets of FOC/96 6349 (Snapshot Search for High Redshift QSOs with Far-UV Flux) L. STIS/CCD 7762 (The Gaseous Environment of the Ursa Minor Dwarf Spheroidal Galaxy) M. Two Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) N. FOC/96 6350 (FOC Spectroscopy of BALQSOs) O. STIS/CCD/MA1 7400 (Evidence from Boron for Massive Star Evolution) P. WF/PC-2 6833 (A Cepheid Search in the Virgo Cluster Galaxy NGC 4571) Q. WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS Galaxy Sample-Part 1 High) R. STIS/CCD/MA1 7480 (The Changing Wind Structure of the Erupting SMC WN/LBV System HD5980) S. WF/PC-2 6363 (The Environments and Host Galaxies of BL-Lac Objects) 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 3 Successful : 3 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: The HST operations team is utilizing the SAMS Back-up facility and the PRS ground system. The spacecraft entered inertial hold safemode at 127/105856Z (HSTAR 7125). The Earth/Moon bright object detection test failed. The spacecraft re-established communications. Operations personnel ensured that the spacecraft was stable and that all subsystems were operating nominally. The Earth/Moon bright object protection test was re-enabled at 127/1854Z. The DCE was powered on at 127/1858Z. Two Health and Safety loads were built by the STScI, uplinked to the spacecraft and executed. They ran from 129/040000Z until 130/000000Z, when normal science commanding was re-established. Per an operations request, the payload recovery COP was completed at 129/0722Z. The NSSC-1 status buffer was dumped at 129/0722Z per ROP NS-3. The spacecraft was recovered from inertial hold at 129/0430Z as directed by an operations request. All science and engineering data was cleared from the SSR at 129/0743Z. Six realtime FHST maps were run throughout the weekend. An attitude reference update was performed at 129/2013Z. Using ROP SR-1, the SSR EDAC error counter was cleared twice. Normal operations commenced at 130/000000Z with the execution of a new science SMS. The first guide star acquisition accurred at 130/0130Z and was successful. Two TTRs were written: one for 4-minute, 5-second recoverable loss of data due to a late acquisition on a TDE event, another for a late acquisition on a TDE event for which the data loss was recoverable. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH